# SketchGuard: Scaling Byzantine-Robust Decentralized Federated Learning via Sketch-Based Screening

### v2

@@ Abstract.
Decentralized Federated Learning (DFL) enables privacy-preserving collaborative training without centralized servers, but remains vulnerable to Byzantine attacks where malicious clients submit corrupted model updates. Existing Byzantine-robust DFL defenses rely on similarity-based neighbor screening that requires every client to exchange and compare complete high-dimensional model vectors with all neighbors in each training round, creating prohibitive communication and computational costs that prevent deployment at web scale. We propose SketchGuard, a general framework that decouples Byzantine filtering from model aggregation through sketch-based neighbor screening. SketchGuard compresses $d$-dimensional models to $k$-dimensional sketches ($k\ll d$) using Count Sketch for similarity comparisons, then selectively fetches full models only from accepted neighbors, reducing per-round communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$, where $|\mathcal{N}_{i}|$ is the neighbor count and $|\mathcal{S}_{i}|\leq|\mathcal{N}_{i}|$ is the accepted neighbor count. We establish rigorous convergence guarantees in both strongly convex and non-convex settings, proving that Count Sketch compression preserves Byzantine resilience with controlled degradation bounds where approximation errors introduce only a $(1+O(\epsilon))$ factor in the effective threshold parameter. Comprehensive experiments across multiple datasets, network topologies, and attack scenarios demonstrate that SketchGuard maintains identical robustness to state-of-the-art methods, with mean test error rate deviation of only up to 0.35 percentage points, while reducing computation time by up to 82% and communication overhead by 50-70% depending on filtering effectiveness, with benefits scaling multiplicatively with model dimensionality and network connectivity. These results establish the viability of sketch-based compression as a fundamental enabler of robust DFL at web scale.
Decentralized Learning, Byzantine Robustness, Dimensionality Reduction, Scalable Machine Learning
copyright: rightsretainedjournalyear: 2026ccs: Security and privacy Systems security
1. Introduction
@@
Federated Learning (FL) enables collaborative training of AI models over distributed data while preserving privacy by keeping raw data local (McMahan et al., 2017). However, the canonical, server-assisted architecture of FL centralizes aggregation of model parameters, creating a single point of failure, a communication bottleneck, and trust issues (Fang et al., 2024). These drawbacks have catalyzed Decentralized Federated Learning (DFL), where clients exchange model updates in a peer-to-peer manner over a dynamic communication graph, thereby improving scalability and resilience (Beltrán et al., 2023).
@@
A central challenge in DFL is byzantine robustness: the ability to withstand malicious clients that submit arbitrary or carefully crafted updates to poison training, induce consensus drift, or trigger targeted failures (Blanchard et al., 2017; Baruch et al., 2019). Unlike centralized FL where robust rules such as Krum (Blanchard et al., 2017), coordinatewise Median (Yin et al., 2018), or Trimmed-Mean (Wang et al., 2025) are applied once at a centralized server, DFL requires every client to aggregate its neighbors’ updates under local, graph-limited views, often with non-IID data and time-varying connectivity. To address this added complexity, most DFL defenses adopt local-consistency filters, where clients accept a neighbor’s update only if it is sufficiently similar to their own state, and then average over the accepted subset (Guo et al., 2021; He et al., 2022; Fang et al., 2024; Cajaraville-Aboy et al., 2024; El-Mhamdi et al., 2021). These mechanisms provide convergence and robustness guarantees in both strongly convex and non-convex model training settings, yet suffer from a fundamental scalability bottleneck: clients must exchange and compare complete, high-dimensional model vectors with all neighbors in every round. For emerging web-scale applications like decentralized training of frontier models with billions of parameters across thousands of distributed participants (Borzunov et al., 2023; Long, 2024), this creates prohibitive communication and computation costs that prevent practical implementation of these systems at scale.
@@
2. Preliminaries and Related Work
@@
2.1. DFL Problem Formulation and Protocol
@@
(1) $\min_{\mathbf{w}\in\mathbb{R}^{d}}F(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{w}),$
@@
(2) $\mathbf{w}_{i}^{t+1/2}=\mathbf{w}_{i}^{t}-\eta\nabla f_{i}(\mathbf{w}_{i}^{t}).$
@@
(3) $\mathbf{w}_{i}^{t+1}=\alpha\mathbf{w}_{i}^{t+1/2}+(1-\alpha)\cdot\mathrm{AGG}_{i}\left(\{\mathbf{w}_{j}^{t+1/2}:j\in\mathcal{N}_{i}\}\right),$
@@
2.2. Byzantine Attack Models
@@
2.3. Byzantine-Robust DFL Defenses
@@
UBAR (Guo et al., 2021) utilizes two-stage neighbor selection based on distance and loss filtering, followed by averaging.
@@
WFAgg (Cajaraville-Aboy et al., 2024) proposes multi-filter approaches for dynamic topologies, combining several screening mechanisms while maintaining the same computational complexity limitations.
However, all these methods require computation of similarities between complete $d$-dimensional model vectors with all neighbors, creating bottlenecks that prevent practical web-scale deployments. Table 1 compares SketchGuard with these existing DFL defense approaches.
Table 1. Comparison of Byzantine-robust decentralized aggregation algorithms. “Convex”/“Non-convex” indicate convergence guarantees; “No $c_{i}$” means no need to know compromised node ratio; “No Complete” means no complete-graph assumption.
Algorithm Convex Non-convex No $c_{i}$ No Complete Scalable
UBAR (Guo et al., 2021) Med.
LEARN (El-Mhamdi et al., 2021) Low
SCCLIP (He et al., 2022) Med.
BALANCE (Fang et al., 2024) Med.
WFAgg (Cajaraville-Aboy et al., 2024) Med.
SketchGuard High
2.4. Compression Techniques in FL
Communication and computational efficiency remain fundamental challenges in FL, motivating extensive research in model compression. Quantization techniques reduce parameter precision (Alistarh et al., 2017; Bernstein et al., 2018), sparsification methods transmit only significant updates (Stich et al., 2018; Sattler et al., 2019), and low-rank approaches exploit model structure (Konečnỳ et al., 2016; Haddadpour et al., 2021). However, most compression techniques target benign settings and often compromise robustness when Byzantine participants are present. Traditional quantization can amplify malicious update impact, while sparsification enables attackers to concentrate influence in transmitted coordinates. More broadly, existing compression methods are incompatible with Byzantine-robust aggregation algorithms, preventing scalable deployment of robust FL.
2.5. Count Sketch
@@
(4) $\mathrm{CS}(\mathbf{w})[b]=\sum_{i:h(i)=b}s(i)w_{i},\quad b=1,\ldots,k.$
@@
Lemma 0 (Distance Preservation (Charikar et al., 2002)).
@@
(5) $(1-\epsilon)\|\mathbf{u}-\mathbf{v}\|^{2}\leq\|\mathrm{CS}(\mathbf{u})-\mathrm{CS}(\mathbf{v})\|^{2}\leq(1+\epsilon)\|\mathbf{u}-\mathbf{v}\|^{2}$
3. SketchGuard: Scalable Robust Aggregation
We now present SketchGuard, which leverages Lemma 1 to perform neighbor screening in the sketch domain. The key design rationale is that if Count Sketch approximately preserves distances, then similarity-based filtering decisions made using sketches will closely match those made using full-precision models, allowing us to defer expensive full-model exchanges until after filtering.
3.1. Protocol Description
Figure 1. The SketchGuard Protocol:
@@ 3.1. Protocol Description
(6) $\|\mathbf{s}_{i}^{t+1/2}-\mathbf{s}_{j}^{t+1/2}\|\leq\gamma\exp(-\kappa t/T)\|\mathbf{s}_{i}^{t+1/2}\|$
@@
Remark 1.
When using Count Sketch compression with approximation parameter $\epsilon$, the effective threshold parameter becomes $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$ to account for distance preservation errors in the sketch domain, as established by the analysis in Lemma 1. This ensures that our theoretical convergence guarantees in Section 4 accurately reflect the impact of compression on filtering decisions.
@@
(7) $\mathbf{w}_{i}^{t+1}=\alpha\mathbf{w}_{i}^{t+1/2}+\frac{(1-\alpha)}{|\mathcal{S}_{i}^{t}|}\sum_{j\in\mathcal{S}_{i}^{t}}\mathbf{w}_{j}^{t+1/2}$
@@
3.2. Complexity Analysis and Performance Trade-offs
@@
Table 2. Computational complexity comparison with state-of-the-art (SOTA) across training phases.
@@
Model Verification & Aggregation $O(d\cdot|\mathcal{N}_{i}|)$ $O(d\cdot|\mathcal{S}_{i}^{t}|)$
Total Per Round $O(d\cdot|\mathcal{N}_{i}|)$ $O(d+k\cdot|\mathcal{N}_{i}|+d\cdot|\mathcal{S}_{i}^{t}|)$
@@ 4.1. Technical Assumptions
Assumption 1 (Strong Convexity).
@@
Assumption 3 (Bounded Variance).
@@
Assumption 4 (Bounded Parameters).
@@
Assumption 5 (Graph Connectivity).
@@
Assumption 6 (Hash Function Synchronization).
All clients use identical hash and sign functions for consistent sketching across the network.
4.2. Main Convergence Results
Leveraging Lemma 1, we establish convergence guarantees for SketchGuard that achieve similar results to those based on full-precision methods.
Theorem 1 (Strongly Convex Convergence with Compression).
@@ 4.2. Main Convergence Results
(8) $\quad+\frac{2L\eta\delta^{2}}{\mu}+\frac{2\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\mu\eta}$
@@
Theorem 2 (Non-Convex Convergence with Compression).
@@
(9) $\quad+4L\eta\delta^{2}+\frac{4\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\eta}$
@@
4.3. Robustness Preservation Under Compression
@@
Lemma 0 (Sketch Compression Preserves Byzantine Resilience).
Under the conditions of Theorems 1 and 2, sketch-based filtering with effective threshold $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$ provides equivalent Byzantine resilience to full-precision filtering with threshold $\gamma_{\text{eff}}$. Specifically, compression does not create new attack vectors beyond the controlled threshold degradation.
@@
This follows directly from the convergence analysis in Theorems 1 and 2. Both proofs bound the neighbor difference terms under sketch-based filtering (Appendix A.5):
@@ 5. Performance Evaluation
We evaluate SketchGuard [^1] Code and experiment artifacts available at doi.org/10.5281/zenodo.17223405. through comprehensive experiments demonstrating that sketch-based compression maintains identical Byzantine robustness to state-of-the-art full-precision defenses while substantially reducing computational and communication overhead.
5.1. Experimental Setup
5.1.1. Datasets and Models
We conduct experiments on two standard federated learning benchmarks from the LEAF suite (Caldas et al., 2018), chosen for their naturally non-IID data distributions that closely reflect real-world heterogeneity in federated environments. We evaluate both robustness and scalability across diverse model sizes.
FEMNIST (Federated Extended MNIST) is a character recognition dataset with 62 classes distributed across 3,550 users, where each user represents a different writer’s handwriting samples. We employ a convolutional neural network comprising two convolutional layers followed by two fully connected layers. Our baseline configuration uses 6.6M parameters, which we systematically scale from 220K to over 60M parameters in our dimensionality scaling experiments (Section 5.3) to evaluate SketchGuard under varying model complexities representative of modern federated learning deployments.
CelebA is a celebrity face attributes dataset distributed across 9,343 users. We use it for binary smile classification (smiling or not smiling), with a CNN architecture adapted for RGB inputs ($84\times 84$ pixels) containing 2.2M parameters. The network consists of two convolutional layers and two fully connected layers. Complete architectural specifications for both models are provided in Appendix C.1.
5.1.2. Network Topologies
(a) Ring(b) ER Sparsep=0.2p=0.2(c) ER Mediump=0.45p=0.45(d) ER Densep=0.6p=0.6(e) Fully Connected
Figure 2. Network topologies used in the robustness evaluation experiments.
We evaluate defense mechanisms across carefully selected network topologies that capture the spectrum of connectivity patterns in real-world decentralized systems, from resource-constrained sparse networks to well-connected infrastructures. For robustness evaluation experiments in Section 5.2, we employ Erdős-Rényi (ER) random graphs (Erdős and Rényi, 1960), which naturally model peer-to-peer networks where connections form probabilistically. ER graphs are generated by independently connecting each pair of nodes with probability $p$, producing networks with variable local connectivity that reflect real-world decentralized learning scenarios where participants discover neighbors organically rather than through engineered topologies. We evaluate five topologies spanning from minimal connectivity (Ring) to maximum connectivity (Fully Connected), with three intermediate ER graph densities as shown in Figure 2; complete specifications are provided in Appendix C.2.1. These experiments use 20 nodes to enable comprehensive evaluation across varying malicious client fractions and defense mechanisms while maintaining computational tractability. For network scaling experiments in Section 5.3, we employ k-regular graphs (Bondy et al., 1976) where each node has exactly $k$ neighbors. This design choice ensures consistent per-node computational load across all defense mechanisms, enabling fair comparison of algorithmic efficiency independent of topology-induced variance. We evaluate networks scaling from 20 to 300 nodes, demonstrating SketchGuard’s scalability to large-scale deployments with hundreds of participants.
5.1.3. Attack Models
We evaluate Byzantine robustness using two representative attack strategies that represent different threat models and attack sophistication levels.
Directed Deviation attacks employ an optimization-based approach that crafts malicious updates by solving for parameters that maximize deviation in the direction opposite to honest gradient descent. This adaptive attack strategy represents a strong adversary that actively attempts to subvert Byzantine-robust aggregation mechanisms by carefully positioning malicious updates to evade detection while maximizing impact. We implement the directed deviation attack following the formulation in (Fang et al., 2020), which has demonstrated effectiveness against common Byzantine-robust aggregation rules including Krum and trimmed mean.
Gaussian attacks (Blanchard et al., 2017) inject random noise $\hat{\mathbf{w}}_{j}=\mathbf{w}_{j}+\mathcal{N}(0,\sigma^{2}\mathbf{I})$ with $\sigma=1$, representing less sophisticated but realistic adversaries that disrupt convergence through stochastic perturbations. While simpler than directed deviation, these attacks model practical scenarios where attackers lack complete knowledge of the aggregation mechanism or resources for optimization-based attacks.
We vary the fraction of Byzantine clients from 0% to 80% in 10% increments to evaluate robustness across different attack intensities. These attack strategies provide complementary evaluation: directed deviation establishes defense performance against sophisticated adaptive adversaries, while Gaussian attacks validate robustness against simpler but more common threat models.
5.1.4. Baselines and Configuration
We compare SketchGuard against four baseline methods: D-FedAvg (Lian et al., 2017), KRUM (Blanchard et al., 2017), UBAR (Guo et al., 2021), and BALANCE (Fang et al., 2024), representing the spectrum from non-robust to state-of-the-art Byzantine-robust aggregation. For all experiments, except the $k$-ablation study, we employ sketch sizes $k=1000$ for FEMNIST and $k=350$ for CelebA. For the baseline model configurations, this achieves compression ratios of 6603:1 and 6342:1 respectively, with approximation error $\epsilon\lesssim 0.2$ and probability $1-\delta>0.99$. All experiments use standard DFL configurations with 10 global rounds and 3 local epochs per round. We repeat each experiment across 3 random seeds and report mean results; observed variances are minimal across all metrics. Complete hyperparameter specifications are provided in Appendix C.2.2.
5.1.5. Evaluation Metrics
We evaluate defense mechanisms using three complementary metrics that capture robustness, computational efficiency, and communication overhead.
Test Error Rate (TER). Defined as 1test accuracy1-$\text${test accuracy}, TER measures performance degradation averaged across honest clients. Lower values indicate better Byzantine resilience.
Per-Round Computation Time. Wall-clock time for neighbor screening and aggregation phases per client per round, excluding local training which remains constant across methods. All measurements are recorded on identical hardware.
Communication Overhead. Total parameters transmitted per client per round, including sketches and full model exchanges. We report both absolute counts and reduction percentages relative to full-precision baselines.
5.2. Byzantine Robustness Evaluation
5.2.1. Attack Strategy Resilience

### v3

@@ Abstract
Decentralized Federated Learning enables privacy-preserving collaborative training without centralized servers but remains vulnerable to Byzantine attacks. Existing defenses require exchanging high-dimensional model vectors with all neighbors each round, creating prohibitive costs at scale. We propose SketchGuard, which decouples Byzantine filtering from aggregation via sketch-based screening. SketchGuard compresses $d$-dimensional models to $k$-dimensional sketches ($k\ll d$) using Count Sketch, then fetches full models only from accepted neighbors, reducing communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$, where $|\mathcal{N}_{i}|$ is the neighbor count and $|\mathcal{S}_{i}|\leq|\mathcal{N}_{i}|$ is the accepted count. We prove convergence in strongly convex and non-convex settings, showing that approximation errors introduce only a $(1+O(\epsilon))$ factor in the effective threshold. Experiments demonstrate SketchGuard maintains state-of-the-art robustness (mean TER deviation $\leq$0.5 percentage points) while reducing computation by up to 82% and communication by 50–70%.
Machine Learning, ICML
1 Introduction
@@
Federated Learning (FL) enables collaborative training of AI models over distributed data while preserving privacy by keeping raw data local (McMahan et al., 2017). However, the canonical, server-assisted architecture of FL centralizes aggregation of model parameters, creating a single point of failure, a communication bottleneck, and trust issues (Fang et al., 2024). These drawbacks have catalyzed Decentralized Federated Learning (DFL), where clients exchange model updates in a peer-to-peer manner over a dynamic communication graph, thereby improving scalability and resilience (Martínez Beltrán et al., 2023).
@@
A central challenge in DFL is byzantine robustness: the ability to withstand malicious clients that submit arbitrary or carefully crafted updates to poison training, induce consensus drift, or trigger targeted failures (Blanchard et al., 2017; Baruch et al., 2019). Unlike centralized FL where robust rules such as Krum (Blanchard et al., 2017), coordinatewise Median (Yin et al., 2018), or Trimmed-Mean (Wang et al., 2025) are applied once at a centralized server, DFL requires every client to aggregate its neighbors’ updates under local, graph-limited views, often with non-IID data and time-varying connectivity. To address this added complexity, most DFL defenses adopt local-consistency filters, where clients accept a neighbor’s update only if it is sufficiently similar to their own state, and then average over the accepted subset (Guo et al., 2022; He et al., 2022; Fang et al., 2024; Cajaraville-Aboy et al., 2025; El-Mhamdi et al., 2021). These mechanisms provide convergence and robustness guarantees in both strongly convex and non-convex model training settings, yet suffer from a fundamental scalability bottleneck: clients must exchange and compare complete, high-dimensional model vectors with all neighbors in every round. For emerging web-scale applications like decentralized training of frontier models with billions of parameters across thousands of distributed participants (Borzunov et al., 2023; Long, 2024), this creates prohibitive communication and computation costs that prevent practical implementation of these systems at scale.
@@
2 Preliminaries and Related Work
@@
2.1 DFL Problem Formulation and Protocol
@@
$\min_{\mathbf{w}\in\mathbb{R}^{d}}F(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{w}),$ (1)
@@
$\mathbf{w}_{i}^{t+1/2}=\mathbf{w}_{i}^{t}-\eta\nabla f_{i}(\mathbf{w}_{i}^{t}).$ (2)
@@
$\mathbf{w}_{i}^{t+1}=\alpha\mathbf{w}_{i}^{t+1/2}+(1-\alpha)\cdot\mathrm{AGG}_{i}\left(\{\mathbf{w}_{j}^{t+1/2}:j\in\mathcal{N}_{i}\}\right),$ (3)
@@
2.2 Byzantine Attack Models
@@
2.3 Byzantine-Robust DFL Defenses
@@
UBAR (Guo et al., 2022) utilizes two-stage neighbor selection based on distance and loss filtering, followed by averaging.
@@
WFAgg (Cajaraville-Aboy et al., 2025) proposes multi-filter approaches for dynamic topologies, combining several screening mechanisms while maintaining the same computational complexity limitations.
However, all these methods require computation of similarities between complete $d$-dimensional model vectors with all neighbors, creating bottlenecks that prevent practical web-scale deployments. A detailed comparison of SketchGuard with existing methods is provided in Appendix C.
2.4 Compression Techniques in FL
Communication and computational efficiency remain fundamental challenges in FL, motivating extensive research in model compression. Quantization techniques reduce parameter precision (Alistarh et al., 2017), sparsification methods transmit only significant updates (Stich et al., 2018; Sattler et al., 2020), and low-rank approaches exploit model structure (Konečný et al., 2016; Haddadpour et al., 2021). However, most compression techniques target benign settings and often compromise robustness when Byzantine participants are present. Traditional quantization can amplify malicious update impact, while sparsification enables attackers to concentrate influence in transmitted coordinates. More broadly, existing compression methods are incompatible with Byzantine-robust aggregation algorithms, preventing scalable deployment of robust FL.
2.5 Count Sketch
@@
$\mathrm{CS}(\mathbf{w})[b]=\sum_{i:h(i)=b}s(i)w_{i},\quad b=1,\ldots,k.$ (4)
@@
Lemma 2.1 (Distance Preservation (Charikar et al., 2002)).
@@
$(1-\epsilon)\|\mathbf{u}-\mathbf{v}\|^{2}\leq\|\mathrm{CS}(\mathbf{u})-\mathrm{CS}(\mathbf{v})\|^{2}\leq(1+\epsilon)\|\mathbf{u}-\mathbf{v}\|^{2}$ (5)
3 SketchGuard: Scalable Robust Aggregation
We now present SketchGuard, which leverages Lemma 2.1 to perform neighbor screening in the sketch domain. The key design rationale is that if Count Sketch approximately preserves distances, then similarity-based filtering decisions made using sketches will closely match those made using full-precision models, allowing us to defer expensive full-model exchanges until after filtering.
3.1 Protocol Description
Figure 1: The SketchGuard Protocol
@@ 3.1 Protocol Description
$\|\mathbf{s}_{i}^{t+1/2}-\mathbf{s}_{j}^{t+1/2}\|\leq\gamma\exp(-\kappa t/T)\|\mathbf{s}_{i}^{t+1/2}\|$ (6)
@@
Remark 3.1.
When using Count Sketch compression with approximation parameter $\epsilon$, the effective threshold parameter becomes $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$ to account for distance preservation errors in the sketch domain, as established by the analysis in Lemma 2.1. This ensures that our theoretical convergence guarantees in Section 4 accurately reflect the impact of compression on filtering decisions.
@@
$\mathbf{w}_{i}^{t+1}=\alpha\mathbf{w}_{i}^{t+1/2}+\frac{(1-\alpha)}{|\mathcal{S}_{i}^{t}|}\sum_{j\in\mathcal{S}_{i}^{t}}\mathbf{w}_{j}^{t+1/2}$ (7)
@@
3.2 Complexity Analysis and Performance Trade-offs
@@
Table 1: Computational complexity comparison with state-of-the-art (SOTA) across training phases.
@@
Model Verification & Aggregation $O(d\cdot|\mathcal{S}_{i}^{t}|)$ $O(d\cdot|\mathcal{S}_{i}^{t}|)$
Total Per Round $O(d+d\cdot|\mathcal{N}_{i}|+d\cdot|\mathcal{S}_{i}^{t}|)$ $O(d+k\cdot|\mathcal{N}_{i}|+d\cdot|\mathcal{S}_{i}^{t}|)$
@@ 4.1 Technical Assumptions
Assumption 4.1 (Strong Convexity).
@@
Assumption 4.3 (Bounded Variance).
@@
Assumption 4.4 (Bounded Parameters).
@@
Assumption 4.5 (Graph Connectivity).
@@
Assumption 4.6 (Hash Function Synchronization).
All clients use identical hash and sign functions for consistent sketching across the network. This is trivially satisfied by seeding the hash functions with the model’s parameter count $d$, which is implicitly shared among all clients training the same architecture.
4.2 Main Convergence Results
Leveraging Lemma 2.1, we establish convergence guarantees for SketchGuard that achieve similar results to those based on full-precision methods.
Theorem 4.7 (Strongly Convex Convergence with Compression).
@@ 4.2 Main Convergence Results
$\quad+\frac{2L\eta\delta^{2}}{\mu}+\frac{2\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\mu\eta}$ (8)
@@
Theorem 4.8 (Non-Convex Convergence with Compression).
@@
$\quad+4L\eta\delta^{2}+\frac{4\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\eta}$ (9)
@@
4.3 Robustness Preservation Under Compression
@@
Lemma 4.9 (Sketch Compression Preserves Byzantine Resilience).
Under the conditions of Theorems 4.7 and 4.8, sketch-based filtering with effective threshold $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$ provides equivalent Byzantine resilience to full-precision filtering with threshold $\gamma_{\text{eff}}$. Specifically, compression does not create new attack vectors beyond the controlled threshold degradation.
@@
This follows directly from the convergence analysis in Theorems 4.7 and 4.8. Both proofs bound the neighbor difference terms under sketch-based filtering (Appendix A.5):
@@ 5 Performance Evaluation
We evaluate SketchGuard [^1] Code and experiment artifacts available at [Anonymized Repository]. through comprehensive experiments demonstrating that sketch-based compression maintains identical Byzantine robustness to state-of-the-art full-precision defenses while substantially reducing computational and communication overhead.
5.1 Experimental Setup
5.1.1 Datasets and Models
We evaluate on three federated learning benchmarks from the LEAF suite (Caldas et al., 2019): FEMNIST (62-class character recognition, 3,550 users, CNN with 6.6M parameters), CelebA (binary smile classification, 9,343 users, CNN with 2.2M parameters), and Sent140 (Twitter sentiment analysis, 660,120 users, LSTM with 1.2M parameters). These datasets provide naturally non-IID distributions reflecting real-world federated heterogeneity across both image and text modalities. Complete architectural specifications are provided in Appendix D.1.
5.1.2 Network Topologies
We evaluate five topologies spanning from minimal connectivity (Ring, degree 2) to maximum connectivity (Fully Connected), with three intermediate Erdős-Rényi densities ($p\in\{0.2,0.45,0.6\}$). The Erdős-Rényi topologies are implemented as dynamic networks where connections are resampled after each round, modeling realistic peer-to-peer environments with changing connectivity. Robustness experiments use 20-node networks; scalability experiments use k-regular graphs scaling from 20 to 300 nodes, ensuring uniform node degree to isolate algorithmic scaling behavior from network topology effects. Complete specifications are in Appendix D.3.1.
5.1.3 Attack Models and Evaluation Metrics
We evaluate four attack strategies: Directed Deviation (Fang et al., 2020), an optimization-based attack maximizing deviation opposite to honest gradients; Gaussian (Blanchard et al., 2017), injecting random noise sampled from the Gaussian distribution; Krum (Fang et al., 2020), exploiting distance-based selection via clustered malicious models; and Backdoor (Bagdasaryan et al., 2020), embedding hidden trigger patterns for targeted misclassification. We vary Byzantine fraction from 0% to 80%. Detailed attack descriptions are in Appendix D.4. We report Test Error Rate (TER) for robustness, Attack Success Rate (ASR) for backdoor resilience, per-round computation time, and communication overhead. The details of these metrics are in Appendix D.2.
5.1.4 Baselines and Configuration
We compare against D-FedAvg (Lian et al., 2017), KRUM (Blanchard et al., 2017), UBAR (Guo et al., 2022), and BALANCE (Fang et al., 2024). We use sketch sizes $k=1000$ (FEMNIST), $k=350$ (CelebA), and $k=180$ (Sent140), maintaining compression ratios of approximately 6600:1 with $\epsilon\lesssim 0.2$. All experiments use 10 global rounds, 3 local epochs, and 3 random seeds. Complete hyperparameters are in Appendix D.3.2.
5.2 Byzantine Robustness Evaluation